OpenStudy (anonymous):

Two letters are randomly picked for the word INFINITESSIMAL. If a letter can be used more than once, calculate the probability that both letters selected are vowels given that the first letter is a vowel.

OpenStudy (anonymous):

probability that first letter picked is a vowel is \[\frac{6}{14}\] so the probability that the next letter picked is a vowel is still \[\frac{6}{14}\] multiply those to get the probability that both letters are vowels, which is \[\frac{9}{49}\approx0.183673469388\]

OpenStudy (anonymous):

the answer is 5/13 lol:)

OpenStudy (anonymous):

5/13? how come?

OpenStudy (anonymous):

because it isn't just asking about the probability of two vowels. its asking the probability of getting two vowels given that the first letter is a vowel.

OpenStudy (anonymous):

you see theres a formula for it

OpenStudy (anonymous):

yes, but a letter can be used more than once.

OpenStudy (anonymous):

yes i know.

OpenStudy (anonymous):

what i did was a bit different

OpenStudy (anonymous):

in that case, the probability would be \[\frac{6}{14}\]

OpenStudy (anonymous):